What is the present value of $ 7 comma 000$7,000 ​received: a. TwentyTwenty years from today when the interest rate is 4 %4% per​ year? b. TenTen years from today when the interest rate is 4 %4% per​ year? c. FiveFive years from today when the interest rate is 4 %4% per​ year?
Answer #1

The present value is calculated using the PV function in excel as follows:

a. =PV(4%,20,0,7000) = 3,194.71

b. =PV(4%,10,0,7000) = 4,728.95

c. =PV(4%,5,0,7000) = 5,753.49